1.
If f(x) = 3x + 5 and g(x) = 2x - 9, find f(x) - g(x).
A. 5x + 4
B. x - 4
C. x + 14
D. 5x - 4
PLEASE HELP

Answer by jim_thompson5910(35256)

f(x) - g(x) = (3x+5) - (2x-9)

f(x) - g(x) = 3x+5 - 2x+9

f(x) - g(x) = x + 14